(1 point) Find y as a function of x if y – 7y + 10y = 12et, y(0) = 10, y(0) = 29, y (0) = 10. y(x) = (21/2)+(41/2)^(2x)-3(1 point) Find a particular solution to y + 36y = –24 sin(6t). yp = 16-3e^(-3t)-8cos(3t)
Answer #1

(1) Given: 9-79 + 109l = 120 Characteristic equation : 83_782 +100 = 0 => 0 (82-70+10) =0 => >= 0, 2,5 Homogeneous Solution:y = c(0) + 222 e 2x + 563e5x + 30% 2x 5 . 4cae & 25 Cze & 3ex (0) = 10 - 10 = ct C2 C3 +3 + C2 + c3 = 7 y (0) = 29 29 = 262